Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

4 letters / 4 envelopes

Expert replies
by yass20015 » Sat Aug 08, 2015 1:47 pm
Tanya prepard 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelope with its correct address. If the 4 letters are to be put into the 4 envelopes at random, what is the probability that only 1 letter will be put into the envelope with its correct address?
Answer: 1/3. Why ? thanks
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Sat Aug 08, 2015 1:56 pm
Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelope with its complete address. If the 4 letters are to be put into the 4 envelopes at random, what is the probability that only 1 letter will be put into the envelope with its correct address ?
A. 1/24
B. 1/8
C. 1/4
D. 1/3
E. 3/8
Let the 4 letters be A, B, C and D.
Total ways to arrange the 4 letters = 4! = 24.
Let the correct ordering of the 4 letters be ABCD.

Write out the ways that ONLY A can be put in the correct position:
ACDB
ADBC
Total ways = 2.

Using the same reasoning, there will be 2 ways that ONLY B can be put in the correct position, 2 ways that ONLY C can be put in the correct position, and 2 ways that ONLY D can be put in the correct position.
Thus, the total number of ways to put EXACTLY 1 letter in the correct position = 2+2+2+2 = 8.

Thus:
P(exactly 1 letter is put in the correct position) = 8/24 = 1/3.

The correct answer is D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by yass20015 » Sun Aug 09, 2015 3:33 am
Thanks Mitch, but I did not get it. Can you please put it with E1,E2,E3,E4 and L1,L2,L3,L4 ? Many thanks!
Join the discussion

by GMATGuruNY » Sun Aug 09, 2015 5:10 am
yass20015 wrote:Thanks Mitch, but I did not get it. Can you please put it with E1,E2,E3,E4 and L1,L2,L3,L4 ? Many thanks!
To my solution above, let's add information about the 4 envelopes.

Let the 4 envelopes be a, b, c and d and the 4 corresponding letters be A, B, C and D.
Let the 4 envelopes be placed on a table in the following order:
a-b-c-d.
Implication:
For every letter to be placed in the correct corresponding envelope, the letters must be arranged in the following order:
ABCD.

Total ways to arrange the 4 letters = 4! = 24.

Write out the ways that ONLY A can be put in the correct position:
ACDB
ADBC
Total ways = 2.

Using the same reasoning, there will be 2 ways that ONLY B can be put in the correct position, 2 ways that ONLY C can be put in the correct position, and 2 ways that ONLY D can be put in the correct position.
Thus, the total number of ways to put EXACTLY 1 letter in the correct position = 2+2+2+2 = 8.

Thus:
P(exactly 1 letter is put in the correct position) = 8/24 = 1/3.

The correct answer is D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by GMATGuruNY » Sun Aug 09, 2015 6:01 am
Alternate approach:

Let the 4 letters be A, B, C and D.

Case 1: Only A is placed in the correct envelope
P(A is placed in the correct envelope) = 1/4. (Of the 4 envelopes, only 1 corresponds to A.)
P(B is placed in a wrong envelope) = 2/3. (Of the 3 remaining envelopes, 2 are wrong.)
Of the 2 remaining letters, at most 1 can be now be placed in the correct envelope.
P(this letter is placed in the WRONG envelope) = 1/2. (Of the 2 remaining envelopes, 1 is wrong.)
P(last letter is placed in a wrong envelope) = 1/1 = 1. (The 1 remaining envelope must be wrong.)
To combine these probabilities, we multiply:
P(only A is placed in the correct envelope) = 1/4 * 2/3 * 1/2 * 1 = 1/12.

Using the same reasoning:
P(only B is placed in the correct envelope) = 1/12.
P(only C is placed in the correct envelope) = 1/12.
P(only D is placed in the correct envelope) = 1/12.

Since any of the probabilities above will yield a favorable outcome, we ADD the fractions:
1/12 + 1/12 + 1/12 + 1/12 = 4/12 = [spoiler]1/3[/spoiler].
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion